# Dynamic Recourse Allocation with Karma: An Experimental Study

### v2

@@ Appendix B Remarks regarding departure from pre-registered analysis plan
Originally, we pre-registered the design of the experiment along with a complete plan for analysis at the Open Science Framework (Elokda and Nax, 2023). The analysis presented in the paper above constitutes a deviation from this pre-registered analysis plan. Here, for completeness, we state the hypotheses that were originally pre-registered, and discuss our motivations for deviating from the pre-registered plan.
@@
    Participants with high urgency spread achieve higher rewards than those with low urgency spread.
As regards the efficiency-related hypotheses (1), (3), (6) and (8), the results of our MWW tests strongly support hypothesis (1), weakly support hypothesis (8), and reject hypotheses (3) and (6). Note, however, that our original plan was to measure realized efficiencies using ex-post means, which was found to be overly sensitive to stochastic effects. Therefore, the analysis presented in Section 4 is based on ex-ante expected medians instead. As regards the fairness-related hypotheses (2), (4) and (7), our fairness analysis supports hypotheses (2) and (7), in particular for the high stake treatments, but not hypothesis (4). These conclusions are drawn based on comparisons of the efficiency gains across deciles, and not as measured by the negated standard deviations of final scores, as we had originally planned, which turned out to be a measure that does not produce interpretable results. Finally, the bidding behaviors portrayed in Figure 4 weakly support hypothesis (5), however, due to the bug in the logging of bids discussed in Section 4.5, we could not rigorously test this hypothesis.
@@ Appendix C Supplementary Results
Table 4 reports the MWW test results when comparing the efficiency gains of the first versus the second halves of the experiments, per treatment. With the exception of the high stake-binary treatment, no statistically significant differences are observed.
Table 4: Mann–Whitney–Wilcoxon (MWW) tests with test statistics $U$ (and associated $p$-values in brackets) of efficiency gains in the first versus second halves of the experiments for all treatments. All treatments have $n_{\textup{karma}}=100$ samples.
@@
5 005.05005.05\,005.05 005.0
($0.9912$)
5 460.55460.55\,460.55 460.5
($0.2610$)
20 878.520878.520\,878.520 878.5
($0.4476$)
@@
4 200.54200.54\,200.54 200.5
($0.0509$)
4 949.54949.54\,949.54 949.5
($0.9028$)
18 357.518357.518\,357.518 357.5
($0.1555$)
@@
18 238.518238.518\,238.518 238.5
($0.1277$)
20 684.020684.020\,684.020 684.0
($0.5544$)
77 729.077729.077\,729.077 729.0
($0.4872$)
Table 5 reports the MWW test results when comparing the efficiency gains of each treatment pair as well as between combined treatments. With the exception of low stake-binary versus high stake-binary (and consequently combined low stake versus high stake), no statistically significant differences are observed.
Table 5: Inter-treatment efficiency gain MWW tests reporting test statistics $U$ (with associated $p$-values in brackets). All treatments have $n=100$ samples.
@@
20 098.520098.520\,098.520 098.5
($0.9324$)

### v3

@@ Appendix B Pre-registered analysis plan
The pre-registration of our experiments included an analysis plan in addition to the design and implementation plan (Elokda and Nax, 2023). The analysis presented in the main paper constitutes a partial deviation from this pre-registered analysis. Here, for completeness, we state the hypotheses that were originally pre-registered, discuss the hypotheses that were addressed by our main analysis, and our motivations for deviating from the analysis of the remaining hypotheses.
The pre-registration reads (Elokda and Nax, 2023, Section 2.1): “We wish to test the following hypotheses. Let us define efficiency as the mean of participant final rewards and fairness as the negated standard deviation of final rewards.
@@
    Participants with high urgency spread achieve higher rewards than those with low urgency spread.”
The group-level efficiency gain, cf. Equation 2, is an affine variant of the mean of participant final rewards aimed at controlling for the random realization of urgency. Accordingly, hypothesis (1) is strongly supported by the MWW tests between karma and random reported in Table 5; hypothesis (6) is rejected by the inter-treatment MWW tests reported in Table 8; and Table 8 also provides weak support for hypothesis (8). Hypothesis (3) is also clearly rejected by investigating the raw group-level efficiency gains reported in Table 4 which all fall below 10% of the full possible efficiency.
As regards hypothesis (5), Figure 5 shows an overall tendency of higher mean bids in high urgency over low urgency. We furthermore performed MWW tests of the bid data in low versus high urgency which showed strongly significantly higher bids in high urgency for all treatments and overall (overall results: low urgency: median $3$, n=11 236n=11\,236; high urgency: median $4$, n=6 706n=6\,706; MWW test U=31 553 792.5U=31\,553\,792.5, $p<0.0001$). For the binary treatments, these tests were conducted using the resulting bids from the binary choice, i.e., zero or half of the present karma. This supports hypothesis (5).
Therefore, the main deviation to the pre-registered analysis plan regards the fairness-related hypotheses (2), (4) and (7). We departed from this analysis because in hindsight we realized that the negated standard deviation is not a good measure of fairness. Namely, it fails to satisfy the basic Pareto principle and leads to uninterpretable results, e.g., a scheme in which everyone performs equally poorly is deemed ‘fairer’ than one where everyone performs better with some variation. Therefore, the fairness analysis performed in Section 4.5 focuses on the notion of Pareto improvements instead. If one accepts this difference in fairness definitions, Figure 4 and Table 9 provide strong support for hypothesis (2); partial support for hypothesis (7), in particular for the high stake treatments; but reject hypothesis (4) since an efficiency-maximizing scheme performs similarly to the theoretical karma Nash (even if it is history-unaware) and leads to Pareto improvements over the experimental karma subjects.
The pre-registration furthermore states (Elokda and Nax, 2023, Section 2.3): “As these are the first karma experiments, we have no benchmarks for expected effect sizes and variances, hence we are unable to run a formal sample size calculator but instead are forced to rely on rules of thumb for sample size and power in order to perform non-parametric tests a la Mann-Whitney-Wilcoxon.” Consistently with this statement, Mann-Whitney-Wilcoxon (MWW) is the main statistical test method used in our analysis.
@@ Appendix C Supplementary Results
C.1 Overview of Earnings
Table 6 shows an overview of the actual total earnings (fixed plus bonus fee) in the experiments.
Table 6: Overview of total earnings ${\phi_{\textup{fix}}}+{\phi_{\textup{bon}}}$ for all treatments. Dropouts for which ${\phi_{\textup{fix}}}={\phi_{\textup{bon}}}=0$ are not included in the statistics. IQR refers to the interquartile range.
@@ C.1 Overview of Earnings
Mean: $\mathdollar 10.03$
Median: $\mathdollar 9.96$
Range: $\mathdollar 3.10-\mathdollar 17.67$
IQR: $\mathdollar 8.59-\mathdollar 11.84$
$\mathdollar 10.14$
$\mathdollar 10.30$
$\mathdollar 1.73-\mathdollar 16.64$
$\mathdollar 8.93-\mathdollar 11.93$
$\mathdollar 10.09$
$\mathdollar 9.96$
$\mathdollar 1.73-\mathdollar 17.67$
$\mathdollar 8.71-\mathdollar 11.84$
@@
Mean: $\mathdollar 9.34$
Median: $\mathdollar 8.92$
Range: $\mathdollar 1.50-\mathdollar 18.10$
IQR: $\mathdollar 7.65-\mathdollar 11.18$
$\mathdollar 8.93$
$\mathdollar 8.99$
$\mathdollar 2.57-\mathdollar 18.38$
$\mathdollar 6.63-\mathdollar 11.01$
$\mathdollar 9.13$
$\mathdollar 8.92$
$\mathdollar 1.50-\mathdollar 18.38$
$\mathdollar 7.23-\mathdollar 11.18$
@@
Mean: $\mathdollar 9.69$
Median: $\mathdollar 9.49$
Range: $\mathdollar 1.50-\mathdollar 18.10$
IQR: $\mathdollar 7.91-\mathdollar 11.58$
$\mathdollar 9.53$
$\mathdollar 9.79$
$\mathdollar 1.73-\mathdollar 18.38$
$\mathdollar 7.45-\mathdollar 11.71$
$\mathdollar 9.61$
$\mathdollar 9.61$
$\mathdollar 1.50-\mathdollar 18.38$
$\mathdollar 7.79-\mathdollar 11.67$
C.2 Robustness Checks: Means vs. Medians; With vs. Without Dropouts
Figure 13 compares the mean group-level efficiency gains, cf. Figure 3(a), to the medians. Means and medians are found to be similar.
(a) Means
(b) Medians
Figure 13: Robustness of Figure 3(a) to the choice of means vs. medians.
Figure 14 tests the robustness of the mean individual-level efficiency gains, cf. Figure 3(b), to both whether medians are analyzed instead of means, and whether dropouts are included or excluded from the analysis. When dropouts are included in the analysis, medians are quantitatively slightly higher than means; while no pronounced differences are observed between means and medians when dropouts are excluded from the analysis. On the other hand, both means and medians are higher when dropouts are excluded from the analysis, and especially so in the lower half of the population, which is consistent with the fact that dropouts perform particularly poorly with persistent zero bids.
(a) Means, with dropouts
(b) Medians, with dropouts
(c) Means, without dropouts
(d) Medians, without dropouts
Figure 14: Robustness of Figure 3(b) to the choice of means vs. medians, and whether dropouts are included in the analysis or not.
Figure 15 performs a similar robustness check as Figure 14, but this time for the fairness analysis of Figure 4. In this finer-grained analysis there are no visible differences between means and medians, while excluding dropouts once again leads to higher aggregate efficiency gains especially in the bottom two deciles. Namely, we observe an almost Pareto improvement over the random benchmark across all deciles when dropouts are excluded.
(a) Means, with dropouts
(b) Medians, with dropouts
(c) Means, without dropouts
(d) Medians, without dropouts
Figure 15: Robustness of Figure 4 to the choice of means vs. medians, and whether dropouts are included in the analysis or not.
C.3 Supplementary Statistical Tests
Table 7 reports MWW test results when comparing the group-level efficiency gains of the first versus the second halves of the experiments, per treatment. No statistically significant differences are observed.
Table 7: Mann-Whitney-Wilcoxon (MWW) tests with test statistics $U$ and associated $p$-values of group-level efficiency gains in the first versus second halves of the experiments, i.e., Equation (2) with scores summed over rounds $t\in\{1,\dots,25\}$ versus $t\in\{26,\dots,50\}$, for all treatments. All treatments have $n=5$ samples.
Richness of scheme
  Binary   Full Range Combined
Urgency process
Low Stake
$U$: $12.0$
pp: $1.0000$
$17.0$
$0.4206$
$59.0$
$0.5205$
High Stake
$U$: $5.0$
pp: $0.1508$
$11.0$
$0.8413$
$32.0$
$0.1859$
Combined
$U$: $37.0$
pp: $0.3447$
$58.0$
$0.5708$
$188.0$
$0.7557$
Table 8 reports MWW test results when comparing the efficiency gains of each treatment pair as well as between combined treatments. With the exception of combined low stake versus high stake, no statistically significant differences are observed.
Table 8: Mann-Whitney-Wilcoxon (MWW) tests reporting test statistics $U$ (with associated $p$-values in brackets) of the inter-treatment group-level efficiency gains, cf. Equation (2). All treatments have $n=5$ samples.
@@ C.3 Supplementary Statistical Tests
$48.0$
($0.9097$)
Table 9 reports MWW test results when comparing the individual-level efficiency gains of karma versus random in each of the lowest three deciles, for each treatment and combined. All higher deciles show similarly significant differences in favor of karma as the third-lowest decile and are omitted for visual clarity. The test results support the findings of Figure 4.
Table 9: Mann-Whitney-Wilcoxon (MWW) tests with test statistics $U$ and associated $p$-values of the individual-level efficiency gains, cf. Equation (1), in each of the three lowest deciles under karma versus random for all treatments. Karma treatments have $n_{\textup{karma}}=10$ samples, and random is based on $n_{\textup{random}}=10\,000$ samples corresponding to simulating random allocation in each experimental group 1 0001\,000 times.
Richness of scheme
Percentile   Binary   Full Range Combined
Urgency process
Low Stake $0-10\%$
$U$: 6 408.56\,408.5
$p$: <0.0001<0.0001
7 315.07\,315.0
<0.0001<0.0001
27 487.027\,487.0
<0.0001<0.0001
$10-20\%$
$U$: 70 351.570\,351.5
pp: $0.0257$
80 424.080\,424.0
$0.0009$
301 552.5301\,552.5
$0.0001$
$20-30\%$
$U$: 99 318.099\,318.0
$p$: <0.0001<0.0001
100 000.0100\,000.0
<0.0001<0.0001
398 933.5398\,933.5
<0.0001<0.0001
High Stake $0-10\%$
$U$: 20 858.020\,858.0
pp: $0.0014$
20 517.520\,517.5
$0.0012$
82 782.582\,782.5
<0.0001<0.0001
$10-20\%$
$U$: 78 054.078\,054.0
pp: $0.0021$
44 359.044\,359.0
$0.5456$
244 131.0244\,131.0
$0.0849$
$20-30\%$
$U$: 100 000.0100\,000.0
$p$: <0.0001<0.0001
98 887.598\,887.5
<0.0001<0.0001
398 130.0398\,130.0
<0.0001<0.0001
Combined $0-10\%$
$U$: 54 084.554\,084.5
$p$: <0.0001<0.0001
52 763.052\,763.0
<0.0001<0.0001
214 243.0214\,243.0
<0.0001<0.0001
$10-20\%$
$U$: 287 382.0287\,382.0
pp: $0.0007$
237983.0 460.5237983.0\,460.5
$0.1404$
1 050 044.51\,050\,044.5
$0.0006$
$20-30\%$
$U$: 399 437.0399\,437.0
$p$: <0.0001<0.0001
5=395 094.55=395\,094.5
<0.0001<0.0001
1 589 324.51\,589\,324.5
<0.0001<0.0001
Table 10 reports MWW test results when comparing the individual-level efficiency gains of karma per each bin of mean absolute differences to Nash portrayed in Figure 7, versus random, for each treatment. The test results support the findings of Figure 7
Table 10: Mann-Whitney-Wilcoxon (MWW) tests with test statistics $U$ and associated $p$-values of the individual-level efficiency gains, cf. Equation (1), per each bin of mean absolute differences to Nash portrayed in Figure 7, versus all simulated random individual samples. Karma bins have varying sample sizes indicated by $n_{\textup{k}}$ in the table, and random is based on $n_{\textup{r}}=100\,000$ samples corresponding to simulating random allocation in each experimental group 1 0001\,000 times.
Mean abs. diff. Richness of scheme
to Nash [karma]   Binary   Full Range
Urgency process
Low Stake 010-1
$n_{\textup{k}}$: 14
$U$: 1 289 244.01\,289\,244.0
$p$: <0.0001<0.0001
22
173 113.0173\,113.0
$0.0733$
121-2
$n_{\textup{k}}$: 55
$U$: 3 981 180.53\,981\,180.5
$p$: <0.0001<0.0001
$27$
2 304 736.02\,304\,736.0
<0.0001<0.0001
232-3
$n_{\textup{k}}$: 21
$U$: 837 073.5837\,073.5
pp: $0.1075$
$50$
3 528 606.03\,528\,606.0
<0.0001<0.0001
343-4
$n_{\textup{k}}$: 4
$U$: 85 666.085\,666.0
pp: $0.0477$
$10$
530 179.5530\,179.5
$0.7410$
454-5
$n_{\textup{k}}$: 4
$U$: 2 767.02\,767.0
pp: $0.0006$
33
46 965.046\,965.0
$0.0393$
>5>5
$n_{\textup{k}}$: 2
$U$: $9.0$
pp: $0.0143$
88
5 074.05\,074.0
<0.0001<0.0001
High Stake 010-1
$n_{\textup{k}}$: $11$
$U$: 1 000 967.51\,000\,967.5
$p$: <0.0001<0.0001
22
199 608.0199\,608.0
$0.0147$
121-2
$n_{\textup{k}}$: 2121
$U$: 1 428 839.01\,428\,839.0
pp: $0.0042$
$14$
1 217 810.51\,217\,810.5
<0.0001<0.0001
232-3
$n_{\textup{k}}$: $59$
$U$: 3 995 844.53\,995\,844.5
$p$: <0.0001<0.0001
$61$
4 272 218.04\,272\,218.0
<0.0001<0.0001
343-4
$n_{\textup{k}}$: 99
$U$: 209 094.5209\,094.5
pp: $0.0054$
2121
541 875.5541\,875.5
$0.0001$
454-5
$n_{\textup{k}}$: 22
$U$: 11 678.011\,678.0
pp: $0.0305$
Table 11 reports MWW test results when comparing the variations in karma and bid distributions of successive rounds, as measured by the Wasserstein-1 distance, between the experimental karma realizations and the simulated stationary Nash benchmark. The test results support the findings of Figure 8.
Table 11: Mann-Whitney-Wilcoxon (MWW) tests with test statistics $U$ and associated $p$-values of the Wasserstein-1 variations in karma and bid distributions between successive rounds, for the experimental MTurk subjects versus the simulated Nash benchmark in all treatments. Karma treatments have $n_{\textup{karma}}=245$ samples (corresponding to 49 rounds in 5 groups), and Nash is based on $n_{\textup{Nash}}=245\,000$ samples corresponding to simulating stationary Nash behavior in each experimental group 1 0001\,000 times.
Karma / Bid Richness of scheme
Distribution   Binary   Full Range
Urgency process
Low Stake Karma
$U$: 21 261 230.021\,261\,230.0
$p$: <0.0001<0.0001
29 481 731.029\,481\,731.0
$0.6318$
Bid
$U$: 21 275 267.021\,275\,267.0
$p$: <0.0001<0.0001
45 444 470.545\,444\,470.5
<0.0001<0.0001
High Stake Karma
$U$: 26 450 120.026\,450\,120.0
pp: $0.0013$
38 284 442.038\,284\,442.0
<0.0001<0.0001
Bid
$U$: 31 678 921.031\,678\,921.0
pp: $0.1324$
53 714 096.553\,714\,096.5
<0.0001<0.0001